OpenStudy (anonymous):

solve for D Q+D=28 .25Q+.1D=5.2

OpenStudy (anonymous):

Q = 28 -D-----let it be equation (1) substitute equation(1) in 25Q + 1D = 5.2 25(28-D) +D = 5.2 i think u can solve for D nw?

OpenStudy (anonymous):

lol k thanks!!

OpenStudy (anonymous):

no probs :)